OpenStudy (anonymous):

Can Someone Help Me find the solution for this equation ? 9 = -2x + 15?

OpenStudy (karatechopper):

9 = -2x + 15 Subtract 15 on both sides 9 = -2x + 15 -15 - 15 ------------- -4=-2x Divide by -2 on both sides... -4=-2x -4/-2=-2x/-2 ?=x I leave that last step for you to do :)

OpenStudy (anonymous):

i got 2 but im pretty sure that may be wrong and too is not even a choice in my thing

OpenStudy (karatechopper):

2 is correct.

OpenStudy (karatechopper):

check to see if you wrote the equation correctly

OpenStudy (anonymous):

no the book is wrong they had 6,12,-12,-21 i wrote the thing right

OpenStudy (karatechopper):

alright. Then I guess the book is wrong :)